### 30 Mar 2026 v2

# RESPOND: Responsive Engagement Strategy for Predictive Orchestration and Dialogue

Abstract.
The majority of voice-based conversational agents still rely on pause-and-respond turn-taking, leaving interactions sounding stiff and robotic. We present RESPOND (Responsive Engagement Strategy for Predictive Orchestration and Dialogue), a framework that brings two staples of human conversation to agents: timely backchannels (“mm-hmm,” “right”) and proactive turn claims that can contribute relevant content before the speaker yields the conversational floor. Built on streaming ASR (Automatic Speech Recognition) and incremental semantics, RESPOND continuously predicts both when and how to interject, enabling fluid, listener-aware dialogue. A defining feature is its designer-facing controllability: two orthogonal dials, Backchannel Intensity (frequency of acknowledgments) and Turn Claim Aggressiveness (depth and assertiveness of early contributions), can be tuned to match the etiquette of contexts ranging from rapid ideation to reflective counseling. By coupling predictive orchestration with explicit control, RESPOND offers a practical path toward conversational agents that adapt their conversational footprint to social expectations, advancing the design of more natural and engaging voice interfaces.
Turn Prediction, Conversational Agents, Voice User Interfaces, Large Language Models, Human-Computer Interaction, User Engagement
conference: Preprint; 2026;
1. Introduction
Voice-based conversational agents have rapidly transitioned from novelty to ubiquity across smartphones, smart speakers, cars, and productivity platforms. Yet their interaction management still diverges from human dialogue: most commercial systems follow a rigid, half-duplex “pause-and-respond” protocol that waits for user silence before producing output. This approach simplifies engineering pipelines but often yields stilted, mechanical exchanges (Skantze, 2021). In contrast, human conversations routinely exhibit sub-second timing, overlaps, and rapid response planning (Levinson and Torreira, 2015). Empirically, shorter response latencies are associated with higher conversational enjoyment, and overlapping interjections do not necessarily reduce perceived quality (Reece et al., 2023). This gap between functional capability and conversational fluency remains a central HCI challenge.
Human conversation is not a sequence of neatly separated turns; it is a continuous, collaborative performance coordinated through subtle verbal and nonverbal cues. Timing, prosody, gaze, and micro-gestures regulate the flow of talk (Levinson and Torreira, 2015). Two mechanisms are especially important for maintaining fluidity: backchannels and cooperative turn claims.
Backchannels (e.g., “mm-hmm,” “right,” “I see”) are short listener acknowledgments produced while the speaker retains the conversational floor. They signal attention and understanding, and their timing is strongly cued by prosodic patterns (Ward and Tsukahara, 2000). When absent in agents, users often attribute the silence to computational delay or a lack of responsiveness. Recent predictive models treat backchannels as interaction opportunities rather than mere reactions, enabling real-time, frame-wise predictions of when and even what to produce (Inoue et al., 2024a, b; Lin et al., 2025).
Cooperative turn claims occur when the listener contributes after or before the speaker has formally yielded the turn, but in a supportive, collaborative manner rather than as a disruptive interruption. Conversation-analytic work shows that some “interruptions” function as cooperative overlap that advances common ground (Weatherall and Edmonds, 2018). In interactive systems, principled barge-in policies and interruption handling reduce latency while avoiding harmful cut-ins (Raux and Eskenazi, 2009; Zhao et al., 2015; Khouzaimi et al., 2018). Moreover, user studies, especially with older adults, report higher perceived naturalness and engagement when agents allow user barge-in and offer timely acknowledgments during speech (Ding et al., 2022; Liu et al., 2025).
In this work, we extend the notion of cooperative turn claims into a unified label, turn claim, representing any listener attempt to claim the conversational floor, whether through interruption, overlap, or smooth turn exchange. This definition contrasts with backchannel, which captures brief, supportive cues that signal engagement without seeking the floor, and stay silent, where the listener provides no verbal response. Representative examples of these three behaviors are illustrated later in Section 3.1.
This paper introduces RESPOND (Responsive Engagement Strategy for Predictive Orchestration and Dialogue), a framework that bridges rigid machine turn-taking and the dynamic qualities of human conversation. RESPOND enables agents to generate both timely backchannels and cooperative turn claims, simulating a full-duplex architecture in which the system listens and speaks concurrently. Built on streaming ASR and incremental semantics, RESPOND continuously analyzes unfolding speech, and its Predictive Orchestration Engine forecasts opportune moments for interjection, mirroring the fluidity of natural dialogue.
A key innovation lies not only in predictive capability but also in explicit controllability. What counts as an appropriate interruption in brainstorming may be intrusive in counseling. RESPOND exposes a designer-facing control paradigm with two orthogonal dials:
  • Backchannel Intensity: tunes the frequency of acknowledgments, from a reserved listener to a highly engaged participant.
  • Turn Claim Aggressiveness: modulates the propensity to contribute proactively, from a system that waits for turn completion to an assertive collaborator.
This level of control is achieved architecturally through lightweight conditioning (e.g., feature-wise modulation), which allows for adjustments without retraining the core model (Perez et al., 2018). By making conversational behavior steerable and transparent, RESPOND aligns with intelligent user interfaces that prioritize user and designer agency.
This work makes the following contributions:
  1. (1)
    We present RESPOND, which integrates predictive models for both backchanneling and cooperative turn claims within a real-time streaming pipeline.
  2. (2)
    We introduce orthogonal dials for Backchannel Intensity and Turn Claim Aggressiveness, providing explicit, fine-grained control over an agent’s conversational style.
  3. (3)
    We benchmark predictive accuracy against state-of-the-art turn-taking/backchanneling models and conduct a pilot study demonstrating the effect of tunable dials on perceived naturalness, responsiveness, and engagement.
Together, these contributions enable conversational agents to respond in ways that feel more collaborative and human-like, bridging the gap between reactive systems and truly interactive dialogue partners.
2. Related Work
2.1. Predictive turn-taking
Classic systems determine speaking opportunities using voice activity detection (VAD) and silence thresholds, often resulting in long pauses and delayed responses. In contrast, modern predictive turn-taking models (PTTMs) aim to forecast both who will speak next and when, allowing agents to plan responses before the conversational floor is released. Voice Activity Projection (VAP) learns to project joint future voice activity directly from audio within short (e.g., 2s) windows, supporting both turn-shift and backchannel inference (Ekstedt and Skantze, 2022b, a). Recent real-time implementations have demonstrated that VAP can operate continuously and efficiently on CPUs (Inoue et al., 2024a). Multimodal extensions such as MM-VAP incorporate facial pose, gaze, and action units to improve hold/shift accuracy, particularly under overlapping speech (Russell and Harte, 2025). Further work has expanded predictive modeling to triadic settings (Lee et al., 2023, 2024; Elmers et al., 2025). Earlier decision-theoretic approaches also framed turn-taking—and even barge-in behavior—as a policy optimization problem (Raux and Eskenazi, 2009; Zhao et al., 2015), while recent surveys synthesize key advances and remaining gaps in the field (Castillo-López et al., 2025; Ji et al., 2024). Yet, existing PTTMs largely fall short in modeling more interactive behaviors such as interruption, overlapping speech, and backchanneling.
2.2. Backchannel and turn claim prediction
Backchannels (e.g., “mm-hmm,” “yeah”) play a critical role in signaling attention and establishing shared understanding. Early efforts to predict backchannel and turn-taking behavior were typically unimodal. For instance, the VAP model (Ekstedt and Skantze, 2022b) mapped acoustic features into a 256-dimensional embedding to estimate the likelihood of turn-keeping, turn-shifting, and backchannel events. In the linguistic domain, Ekstedt and Skantze (Ekstedt and Skantze, 2020) fine-tuned GPT-2 (Radford et al., 2019) with explicit TURN tokens, achieving promising performance. More recent studies (Shukuri et al., 2023; Kim et al., 2025) highlight that large language models (LLMs) are particularly effective at extracting contextual cues from text, often outperforming conventional approaches. Beyond speech and text, visual cues such as eyebrow and mouth movements have also been shown to shape turn-taking in face-to-face interactions (Lee and Narayanan, 2010).
Building on these foundations, multimodal approaches integrate complementary signals to improve robustness. Chang et al. (Chang et al., 2022) classified turn-taking into six sub-categories and introduced an end-to-end ASR-based framework that fused acoustic and linguistic features for both turn-taking and backchannel prediction. Yang et al. (Yang et al., 2022) further demonstrated that gated fusion of acoustic and semantic inputs improved accuracy, especially when textual cues augmented audio-only models. Other works (Kurata et al., 2023; Wang et al., 2024) have similarly combined multimodal signals, achieving state-of-the-art results, while extensions to triadic interactions further highlight the benefits of cross-modal fusion (Lee and Deng, 2024).
Despite these advances, unimodal systems powered by the latest LLMs remain attractive for real-time deployment as they can achieve competitive accuracy while offering lower latency and reduced computational overhead compared to full multimodal pipelines.
2.3. Proactive, barge-in–capable voice agents
User studies consistently report that agents feel more natural when they acknowledge, tolerate, and perform interactions appropriately in flight. For older adults, TalkTive showed that agent backchannels improve engagement during sensitive screening dialogs (Ding et al., 2022). Follow-ups that support both interruptions and backchannels find higher perceived naturalness, fluency, and engagement than strict turn-by-turn baselines (Liu et al., 2025). Beyond lab settings, meeting-scale work has found failed interruptions to improve inclusiveness in remote collaboration (Fu et al., 2022). Linguistic and HCI studies highlight that “interruptions” are partly cultural and can be cooperative overlap, not just dominance (Hilton, 2018, 2016; Weatherall and Edmonds, 2018; Cantrell, 2013).
Beyond speech, overlap also benefits text-based agents. Allowing controlled “textual over-talk” (parallelized messages) improves throughput and perceived flow, echoing speech literature on overlap as a cooperative device (Kim et al., 2025).
Our prototype integrates streaming ASR with incremental semantics extracted from an LLM, enabling the RESPOND module to determine in real time whether the agent should respond.
3. Methodology
We developed a text-based module for predicting listener behaviors such as backchannels and interruptions in real-time conversation. The system takes transcribed speech from a voice-based conversational agent and continuously outputs predictions for when the agent should backchannel, attempt to take the floor, or stay silent. Our design emphasizes lightweight modeling for low-latency deployment and controllability through scalar parameters that modulate the frequency of backchannels and the aggressiveness of interruptions.
3.1. Data Preparation and Analysis
We use two text-based conversational corpora annotated for turn-taking behaviors: the MM-F2F dataset (Lin et al., 2025) and the CANDOR corpus (Reece et al., 2023).
Table 1. Examples of input utterances and their corresponding labels from the MM-F2F dataset.
Input Label
last week yes we stayed home and we KEEP
is hard to decide i just can’t pick one BACKCHANNEL
hi hi good afternoon how are you doing TURN
The MM-F2F dataset consists of face-to-face conversations with multimodal streams, from which we use only the transcripts. For the text-only task and for comparison with prior work, this dataset offers a train-ready format where incoming utterances are annotated with listener response types: TURN, BACKCHANNEL, and KEEP. However, the paper overlooks interruption and overlap behaviors. Table 1 presents three illustrative examples, each corresponding to a different label.
The CANDOR corpus (Reece et al., 2023) comprises over 850 hours of video-call conversations with detailed annotations. Unlike MM-F2F, this dataset requires additional collation and preprocessing before it can be used for training. We first derive frame-level speaking status by aligning word-level transcript timestamps, and we adopt the Backbiter [^1] Backbiter is a turn-model introduced in Reece et al. (2023) that identifies short listener utterances, typically fewer than three words, composed largely of backchannel lexicon items, and not beginning with self-referential phrases (e.g., “I’m…”), as backchannels that occur in parallel with the speaker’s turn rather than as interruptions. format to extract explicit backchannel instances as marked in the corpus.
The transcripts were segmented into 5-second windows with a 50 ms stride, yielding near–word-level inference. In rare cases of very rapid speech (e.g., “in the”), multiple words may fall within the same 50 ms interval. Each window was represented from two perspectives: $w^{A}$, where participant A is treated as the listener and participant B as the speaker, and $w^{B}$, where A and B switch roles. From each perspective, the model is trained to predict the appropriate class at the next word boundary given the listener’s viewpoint. This dual-perspective framing ensures that every conversation contributes two complementary sets of training data—one for A-as-listener and one for B-as-listener. Combined with the fine temporal resolution, it allows the model to capture rapid backchannels or short interruptions multiple times across overlapping windows, providing dense and balanced training coverage.
Figure 1. Illustration of the three annotation labels—turn claim, backchannel, and stay silent—and how they appear in different conversational contexts. Left: listener actions categorized as turn claims (interruption, overlap, turn-taking). Right: corresponding backchannel examples occurring in similar contexts. Gray-shaded regions represent stay silent intervals where the listener produces no response. Color bars indicate who is speaking or backchanneling.
All label situations
Following the conceptual definition introduced in Section 1, we then operationalize three listener behaviors for modeling: turn claim, backchannel, and stay silent. Figure 1 illustrates representative situations for each category. We define turn claim as any listener attempt to take the conversational floor, not necessarily resulting in a turn shift. Such attempts may occur through (i) interruption, where the listener barges in and halts the speaker; (ii) overlapping speech, where the listener speaks simultaneously with the speaker; or (iii) normal turn-taking, where the listener takes the floor after the speaker finishes. In contrast, backchannel refers to the same timing contexts, but the listener produces brief supportive cues (e.g., “uh-huh,” “yeah”) without attempting to claim the floor, following the annotation scheme of the corpus. Finally, any segment in which only the speaker talked—without interruption, overlap, or backchannel from the listener—was labeled as stay silent. Each word boundary was annotated with one of these categories by aligning transcript content with the corresponding timing metadata.
Figure 2. Label distributions before and after downsampling.
Showing how we deal with imbalanced data.
To address class imbalance and reduce biases from variable word-count within a window, we applied bin-stratified balanced downsampling. Figure 2 illustrates the effect of this procedure. The leftmost histogram shows the original distribution of window lengths, which is highly skewed. If we instead perform global class-balanced downsampling (middle), label proportions are equalized overall, but short windows of one or two words remain dominated by the turn claim class. This occurs because many backchannel windows consist of brief acknowledgments such as “yeah” or “interesting,” which are often followed by a speaker response and thus labeled as turn claim. Our method (right) alleviates this issue by balancing labels separately within each length bin.
Concretely, for each window $w_{i}$ we first computed its word-count $C_{i}=\bigl|\text{split}(w_{i})\bigr|$. We then assigned windows to predefined bins: $\{1\},\{2\},\;\{3\text{--}4\},\;\{5\text{--}6\},\;\ldots,\;\{29\text{--}30\},\;\{31+\}$. The first two bins are treated separately because single-word and two-word windows exhibit distinct dynamics: most backchannels fall within these lengths, while single-word windows are often dominated by turn claim responses. Grouping them together would therefore preserve a residual imbalance just like the global class-balanced downsampling in the middle of Figure 2. Within each bin $b$, we considered all labels $y\in\mathcal{Y}$ (with $\mathcal{Y}=\{\textit{turn claim},\textit{backchannel},\textit{stay silent}\}$) and determined the smallest class count in that bin:
(1) $n_{b}=\min_{y\in\mathcal{Y}}\;\#\{\,~(w_{i},y_{i})\mid C_{i}\in b,\;y_{i}=y\,\}$
We then randomly sampled $n_{b}$ windows for each label, ensuring that every $(\text{bin},\text{label})$ bucket contained the same number of examples. Before downsampling, the dataset contained 132,277 turn claim, 159,787 backchannel, and 4,277,563 stay silent windows, showing a strong class imbalance. After applying the bin-stratified balanced downsampling, each class was reduced to 127,674 samples, resulting in a balanced dataset. The downsampled sets were then concatenated, shuffled, and split into training, validation, and test subsets with a ratio of 18:1:1.
3.2. Model
The backbone of our system is the Qwen3-0.6B language model (Qwen et al., 2025), a 0.6B-parameter Transformer designed for efficiency. We fine-tune the model using Low-Rank Adaptation (LoRA (Hu et al., 2022)) to adapt it for turn-taking prediction while keeping the parameter footprint manageable. The model outputs a classification over three categories: turn claim, backchannel, or stay silent.
3.3. Controllability
A key goal of our framework is to allow flexible control over the conversational style of the agent. To this end, we introduce two continuous control parameters: backchannel intensity ($c_{\text{bc}}$) and turn claim aggressiveness ($c_{\text{tc}}$). These values capture how frequently a listener tends to produce backchannels or take the floor during a conversation. By conditioning the model on these values, we can bias predictions toward more passive or more assertive behaviors without retraining the underlying network.
Calculation of control parameters.
The values for the control parameters are computed at the conversation level for each participant. In a dyadic conversation, this yields two distinct sets of control parameters, one for participant A and one for participant B.
Suppose a conversation consists of $N$ total frames. Let $N_{\text{bc}}^{A}$ and $N_{\text{spk}}^{A}$ denote the number of frames in which participant A is backchanneling and speaking, respectively, and let $N_{\text{bc}}^{B}$ and $N_{\text{spk}}^{B}$ denote the corresponding counts for participant B. The control parameters are then defined as:
$c_{\text{bc}}^{A}=\frac{N_{\text{bc}}^{A}}{N},\qquad c_{\text{tc}}^{A}=\frac{N_{\text{spk}}^{A}}{N},\qquad c_{\text{bc}}^{B}=\frac{N_{\text{bc}}^{B}}{N},\qquad c_{\text{tc}}^{B}=\frac{N_{\text{spk}}^{B}}{N}.$
Here, $c_{\text{bc}}$ reflects the proportion of time a participant produces backchannels, while $c_{\text{tc}}$ represents their overall turn claim ratio—that is, the proportion of frames in which the participant speaks (voiced activity excluding backchannels), regardless of whether they currently hold the conversational floor. These counts are computed over the entire conversation, capturing each participant’s general speaking style rather than properties of individual windows. When extracting training segments, each window inherits the control values of the designated listener. As a result, the same conversation yields two complementary sets of windows: one where A is modeled as the listener (with $(c_{\text{bc}}^{A},c_{\text{tc}}^{A})$), and another where B is modeled as the listener (with $(c_{\text{bc}}^{B},c_{\text{tc}}^{B})$). This ensures that control parameters remain consistent with the overall behavioral tendencies of the selected listener.
Figure 3. Mapping control parameters to a uniform scale. Here, $c_{\text{bc}}$ denotes backchannel intensity and $c_{\text{tc}}$ denotes turn claim aggressiveness. The left column shows the raw distributions of the calculated ratios, the middle column shows the distributions after quantile-based transformation, and the right column visualizes the scaled values as both scatter points and a heatmap.
Shows how we scale the control values and they are uniformly distributed.
Instead of using the raw control parameters directly, we normalize them with a quantile-based transformation to distribute the values more evenly over the range $[0,1]$. Specifically, we apply the QuantileTransformer from scikit-learn (Pedregosa et al., 2011), which maps the empirical distribution of the data to a uniform distribution such that each quantile is equally represented. Figure 3 illustrates the distributions before and after transformation. This step addresses the skewed nature of the raw ratios: for example, many conversations have backchannel intensities $c_{\text{bc}}$ concentrated below $0.15$. Without rescaling, most values would lie in a narrow band near zero, making it difficult for the model to learn meaningful distinctions. By contrast, the quantile transformation spreads these differences across the full $[0,1]$ range, improving training stability. In addition, the rescaled values provide a more intuitive interface for controllability at inference time, allowing users to interpret settings on a normalized 0–100% scale rather than in dataset-specific raw proportions.
Integration into the model.
We condition the model on these control values using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018; Dumoulin et al., 2018). Given a hidden representation $h\in\mathbb{R}^{d}$ from the backbone encoder, we compute:
$\text{FiLM}(h;c)=\gamma(c)\odot h+\beta(c),$
where $c=[c_{\text{bc}},c_{\text{tc}}]$ and $\gamma(\cdot),\beta(\cdot)$ are small multi-layer perceptrons that transform the control scalars into per-dimension scaling and shifting factors. This operation adjusts the hidden activations in a way that reflects the overall interaction style of the conversation, without modifying the pretrained backbone. During training, each input window is paired with the $(c_{\text{bc}},c_{\text{tc}})$ values computed from its parent conversation. This ensures that the model learns not only local turn-taking dynamics, but also how these dynamics should adapt under different global styles. For example, in a conversation with a high $c_{\text{bc}}$, the model sees all windows conditioned on a “frequent backchannel” setting, encouraging predictions that align with that style.
Figure 4. System pipeline of our conversational agent module.
System logic and flow.
3.4. System Design
Figure 4 illustrates the end-to-end inference pipeline of RESPOND. Incoming speech is first transcribed into word-level transcripts by an automatic speech recognition (ASR) module. The ASR is powered by Microsoft Azure Speech Cognitive Services. [^2] https://speech.microsoft.com/portal The transcripts are accumulated into sliding windows of up to 5 seconds for inference.
This pipeline supports real-time backchannel and turn claim prediction based on the most recent partial transcript.
Although inference itself operates at word-level granularity, overall responsiveness is constrained by ASR latency—the time between a spoken word and its textual availability. In our implementation, this delay typically ranges between 250–500 ms under stable network conditions, consistent with modern streaming ASR benchmarks. Consequently, the timing of generated backchannels may be offset by this recognition latency, an inherent limitation of text-based real-time systems.
RESPOND module provides user-level customization of interaction style by exposing controllable parameters. In particular, users can specify $(c_{\text{bc}},c_{\text{tc}})$ to adjust the agent’s communication style: higher $c_{\text{bc}}$ encourages more frequent backchannels, while higher $c_{\text{tc}}$ increases the likelihood of overtaking the conversational floor:
  • Passive listener: $c_{\text{bc}}=0.1$, $c_{\text{tc}}=0.0$ (rare backchannels, never interrupts).
  • Collaborative listener: $c_{\text{bc}}=0.6$, $c_{\text{tc}}=0.2$ (frequent backchannels, occasional turn-taking).
  • Assertive speaker: $c_{\text{bc}}=0.1$, $c_{\text{tc}}=0.8$ (infrequent backchannels, frequent interruptions or floor-taking).
Alternatively, the system can estimate $c_{\text{bc}}$ and $c_{\text{tc}}$ dynamically from past conversation history, allowing the agent to mirror the style of its partner. In either case, controllability enables fine-grained adaptation of behavior with negligible computational cost. Then, the resulting decision is passed to a text-to-speech (TTS) system that generates the corresponding verbal response. This architecture ensures that the agent can provide timely acknowledgments or strategically take the floor, creating a more natural interaction flow.
3.5. Experiments
Implementation Details
All experiments were implemented in PyTorch and trained on a single NVIDIA A100 GPU. We used the Qwen3-0.6B language model as the backbone, which has a hidden size of 768. For controllability, we injected two scalar values, backchannel intensity $c_{\text{bc}}$ and turn claim aggressiveness $c_{\text{tc}}$, through a FiLM layer. Each FiLM layer consisted of two one-layer MLPs with ReLU activations, which projected the control scalars into scaling and shifting vectors of size 768 to match the backbone hidden states.
Models were fine-tuned using LoRA with rank $r=16$ and $\alpha=32$, enabling efficient adaptation with a limited parameter budget. Training was performed with a per-device batch size of 128 for both training and evaluation. We optimized using AdamW with a learning rate of $1\times 10^{-5}$, weight decay of 0.01, cosine learning rate scheduling, and a warm-up ratio of 0.1. Gradient clipping was applied with a maximum norm of 1.0. We monitored validation loss at the end of each epoch and selected the checkpoint with the lowest loss as the final model. Training was terminated after three epochs, as validation performance stabilized and earlier checkpoints showed no further improvement. These settings provided stable convergence and good generalization while keeping computational time feasible for rapid iteration. A complete training run required approximately 1.5 hours on the A100 GPU. We fixed the random seed to 42 and disabled data loader multiprocessing to ensure reproducibility.
3.6. Pilot Study
To complement the quantitative evaluation, we conducted a small-scale exploratory pilot study to examine how users perceived and interacted with the controllability features of our conversational agent. Six participants (4 female, 2 male) were recruited from our lab community; all had prior experience with commercial voice assistants but no prior exposure to our system. Figure 7 shows the user interface, which provides two sliders that allow users to adjust backchannel intensity and turn claim aggressiveness on a continuous scale from 0 to 1.
Procedure
The study was reviewed and approved by the appropriate ethics review process at the authors’ institution and was determined to involve minimal risk. No personally identifying data were collected. Each participant first engaged in a 10–12 minute open-ended conversation with the system under a default configuration. Participants were instructed to interact naturally, as they would with a voice-based assistant, by asking questions, sharing opinions, or describing daily experiences. The goal was to elicit spontaneous, mixed-initiative exchanges rather than task-specific or scripted interactions. After this initial interaction, participants were introduced to the controllability interface, which exposed the two scalar parameters $(c_{\text{bc}},c_{\text{tc}})$. They were encouraged to experiment with different settings by adjusting the sliders and immediately observing the changes in the system’s behavior. Once they were comfortable, participants were asked to continue conversing with the agent while iteratively tuning the values until they arrived at what they considered their “ideal” configuration. After the session, participants completed a short semi-structured interview focusing on naturalness, responsiveness, and overall satisfaction with both the default and self-selected settings.
4. Results
4.1. Benchmarking Against Prior Work
We compare RESPOND against previously reported baselines from the MM-F2F benchmark (Lin et al., 2025), which re-implemented and evaluated several representative models on the same dataset: (1) TurnGPT (Ekstedt and Skantze, 2020), which processes only text and predicts turn-hold versus turn-shift actions; (2) the model of Wang et al. (2024), which integrates text and audio features for joint turn-taking and backchannel prediction; and (3) the multimodal model of Kurata et al. (2023), which fuses text, audio, and video streams for turn-taking prediction. For completeness, we also include the official MM-F2F baseline results (both multimodal and text-only configurations) reported by Lin et al. (2025), and add our own results obtained using the same data splits and evaluation protocol.
As shown in Table 2, RESPOND outperforms the text-only approaches, achieving higher F1-scores on the Keep and Turn classes as well as improved overall accuracy compared to the strongest text-only baseline. Although our framework does not surpass the latest multimodal system that combines text, audio, and video, it achieves competitive performance while being substantially more lightweight and efficient—making it suitable for real-time deployment.
All MM-F2F results are reported on the dataset’s native label distribution, without additional class rebalancing. Control parameters for backchannel intensity ($c_{\text{bc}}$) and turn claim aggressiveness ($c_{\text{tc}}$) are not applied in this experiment, as the MM-F2F corpus does not provide annotations for interruptions or overlapping speech. These controllable parameters are introduced and analyzed later using the CANDOR corpus, which explicitly includes such phenomena.
Table 2. The comparison study results on the MM-F2F dataset. We compare TurnGPT, Wang et al., Kurata et al., and the work from MM-F2F (Lin et al., 2025) itself to our model. The T, A, V denote input modalities of text, audio, and video.
Method Modal Acc. F1-Score
Keep Turn BC
TurnGPT (Ekstedt and Skantze, 2020) T 0.645 0.745 0.420 -
Wang et al.’s T+A 0.737 0.742 0.739 0.680
Kurata et al.’s T+A+V 0.720 0.729 0.728 0.667
MM-F2F (Lin et al., 2025) T+A+V 0.823 0.806 0.811 0.906
$M$-F2F T 0.751 0.747 0.767 0.707
Ours T 0.756 0.770 0.773 0.697
4.2. Performance on the CANDOR Corpus
We further evaluated our framework on the CANDOR corpus (Reece et al., 2023), which comprises over 850 hours of video-call conversations annotated with detailed turn-taking behaviors. Unlike MM-F2F, which focuses on clean two-party exchanges, CANDOR captures a richer range of spontaneous conversational phenomena, including interruptions, overlaps, and frequent backchanneling, making it a challenging and comprehensive benchmark for testing both prediction accuracy and controllability.
For this experiment, we used the balanced training split provided by CANDOR to ensure equal representation of turn claim, backchannel, and stay-silent behaviors. We also incorporated the controllability parameters introduced in Section 3.3: $c_{\text{bc}}\in[0,1]$ for Backchannel Intensity and $c_{\text{tc}}\in[0,1]$ for Turn Claim Aggressiveness. These scalar values were concatenated with the model’s hidden representation through FiLM conditioning, allowing fine-grained modulation of listener behavior during both training and inference.
Table 3 summarizes the model’s performance on the CANDOR test set. The model achieved F1-scores of 0.847 for turn claim, 0.845 for backchannel, and 0.902 for stay-silent, with an overall accuracy of 0.87. Figure 5 shows the confusion matrix, illustrating that most misclassifications occur between stay silent and backchannel. This confusion likely arises from conversational contexts in which a brief backchannel could be appropriate but remaining silent is also socially acceptable.
We report only our model’s results on this corpus, as existing baselines such as TurnGPT or the MM-F2F implementations were not designed to handle interruption or overlap categories and lack controllable inputs for backchannel and turn claim behaviors. Moreover, these models were originally optimized for dyadic turn-shift prediction under fixed label schemes, making direct retraining or comparison on CANDOR non-trivial and outside the scope of our controllability study. The purpose of this evaluation is therefore to assess the generalization and tunability of our controllable framework in a complex, naturalistic conversational setting.
Table 3. Performance of our model on the CANDOR corpus.
Category Precision Recall F1
Turn claim 0.8657 0.8290 0.8470
Backchannel 0.8286 0.8614 0.8447
Stay silent 0.8975 0.9062 0.9018
Overall Accuracy 0.87
Figure 5. Confusion matrix on the CANDOR test set.
confusion matrix of the result
4.3. Case Studies
Figure 6. Illustrative examples of RESPOND predictions across three scenarios: a simple greeting, a task-oriented exchange, and a longer casual conversation.
Show word by word prediction for different cases.
In addition to quantitative evaluation with accuracy and F1-scores, we conducted a qualitative analysis to better understand the model’s real-time behavior. For this purpose, we randomly selected representative conversation segments, ranging from simple greetings to task-oriented and free-form interactions, and visualized how prediction probabilities evolved as words were incrementally appended to the observation window. These case studies reveal how the model differentiates among backchannel, turn claim, and stay-silent behaviors, illustrating its sensitivity to conversational context and timing.
Figure 6(a) presents a simple greeting. The model predicts a turn claim slightly early right after the word “you,” though this can be considered reasonable since conversational overlaps often occur immediately after phrases such as “how are you.” Figure 6(b) illustrates a task-oriented interaction where the speaker provides a phone number. Although the model incorrectly predicts a turn claim at the second word, it successfully identifies the natural boundary at the end of the phone number and appears to capture the expected length of a formal phone number. Figure 6(c) shows a longer, casual conversation with hesitations. In this example, the model produces timely backchannels that encourage the speaker to continue, demonstrating context-sensitive responsiveness.
4.4. User Experience
Figure 7. The user interface of our system.
UI
Findings
Participants unanimously reported that adjusting the controllability parameters led to noticeable and meaningful differences in the system’s conversational style. Higher $c_{\text{tc}}$ values were described as making the agent “too pushy” or “constantly jumping in,” while lower $c_{\text{tc}}$ produced a calmer, more comfortable interaction. Backchannel intensity ($c_{\text{bc}}$) was generally well received when moderate, with participants noting that it “felt like someone was listening” or “kept the conversation flowing.” However, excessive backchannels, especially mid-sentence ones, were considered distracting and unnatural.
Figure 8. Distribution of user-selected control parameters.
user control parameters distribution
Preferences
Figure 8 shows the distribution of user-selected control parameters. Three out of six participants ultimately set $c_{\text{tc}}$ close to 0.2, indicating a preference for minimal turn claim. At the same time, most participants selected a moderate $c_{\text{bc}}$, suggesting that occasional, well-timed backchannels are desirable. Two participants opted for slightly higher backchannel intensities, stating that it made the system feel more engaged, but even they avoided high overtaking aggressiveness. Across the group, participants emphasized that the “ideal” configuration balanced attentiveness with restraint.
From both the conversations and the interviews, several key observations emerged:
  • Timing matters: sentence-final backchannels felt natural, but mid-sentence backchannels were disruptive.
  • Overtaking caution: frequent or mistimed turn claims eroded trust and willingness to engage.
  • Perceived control: participants appreciated being able to adjust values themselves, describing the sliders as “knobs” to personalize the system’s personality.
  • Exploration benefit: letting users experiment helped them understand the range of behaviors and settle on comfortable defaults.
Overall, the study confirmed that the controllability framework produces clear, perceivable effects on conversational style, and that users prefer to operate at low overtaking aggressiveness with moderate backchannel intensity. The ability to interactively adjust these settings was particularly valued, as it gave participants a sense of agency and personalization. These findings suggest that controllability is not only technically effective but also user-friendly, though future work is needed to refine backchannel timing and reduce overly aggressive floor-taking.
5. Discussion
RESPOND operationalizes two long-standing desiderata for conversational agents, timely backchannels and cooperative, context-appropriate turn claims, through a text-only, low-latency module with explicit behavioral control. Across MM-F2F and CANDOR, a lightweight unimodal approach delivers competitive accuracy while exposing interpretable dials that practitioners can tune to the social norms of a setting (e.g., ideation vs. counseling). On MM-F2F, we show parity or modest gains versus text-only baselines while remaining far simpler than multimodal stacks; on CANDOR, the model maintains balanced performance across turn claim, backchannel, and stay-silent despite naturalistic overlaps and interruptions.
Our exploratory pilot user study suggests that this steerability is practically useful. Most participants preferred low turn claim aggressiveness ($c_{\text{tc}}\approx 0.2-0.4$) with moderate backchannel intensity ($c_{\text{bc}}\approx 0.4-0.6$), highlighting the importance of sentence-final timing for acknowledgments. A key design implication is that systems should default to conservative settings while allowing for user-specific adjustments.
RESPOND complements emerging ASR–LLM–TTS pipelines that already excel at streaming recognition and generation. Within these pipelines, RESPOND provides policy-level guidance on when and how to interject. Its FiLM-based conditioning integrates directly with a text encoder and can be updated online without retraining. Moreover, the system exposes quantile-scaled control parameters ($c_{\text{bc}}$, $c_{\text{tc}}$), which act as intuitive, human-interpretable controls, the same UI “knobs” that end users adjusted in our pilot study, highlighting a feasible path toward personalized, etiquette-aware conversational agents in production.
6. Limitations and Future Work
Our current model is text-only; while this design maximizes deployability and minimizes latency, it omits prosodic and visual cues that are well known to signal backchannels and turn transitions. Future work may consider extending the model with prosody-aware features (e.g., pause length, pitch movement) and lightweight visual signals (e.g., head nods, eyebrow raises), incorporated through late fusion to preserve low runtime costs. For training, we currently collapse normal turn-taking, interruption, and overlap into a single turn claim class, simplifying dataset unification but masking socially important distinctions. A promising direction is to separate these subtypes and examine how controllability dials differentially modulate them—for instance, encouraging cooperative overlap while suppressing disruptive interruption.
[…]